Calculate the solubility of Ni(OH)2 in pure water and in a 0.0090M NaOH solution. Given that Ksp = (5.48 x 10^-6).

Round answers to 2 significant figures, and have them in g/L units.
Answer #1

NiloH), (s) ĭ Nil (4) + 2014 (4) [Nje t] =5 [oH]=25 (two hydroxide ions are there) Ksp = (s) (95)2 43 = 5.48x10 s3 = 1.37x106